The two cars started a race with 6 and 9 speeds. Their acceleration is 0.7 and 0.5. Respectively What is the distance traveled by the first car when the two cars reach the other end at the same time?
1
Expert's answer
2020-11-27T00:40:30-0500
"s_1=s_2\to v_1t+0.5a_1t^2=v_2t+0.5a_2t^2\\6t+0.5(0.7)t^2=9t+0.5(0.5)t^2\\t=30\ s\\s_1=6(30)+0.5(0.7)(30)^2=495\ m"